Financial data for Joel de Paris, Inc., for last year follow:

Joel de Paris, Inc.
Balance Sheet
Beginning
Balance
Ending
Balance
Assets
Cash $ 133,000 $ 139,000
Accounts receivable 333,000 487,000
Inventory 574,000 484,000
Plant and equipment, net 840,000 810,000
Investment in Buisson, S.A. 407,000 428,000
Land (undeveloped) 254,000 254,000
Total assets $ 2,541,000 $ 2,602,000
Liabilities and Stockholders' Equity
Accounts payable $ 377,000 $ 346,000
Long-term debt 1,041,000 1,041,000
Stockholders' equity 1,123,000 1,215,000
Total liabilities and stockholders' equity $ 2,541,000 $ 2,602,000
Joel de Paris, Inc.
Income Statement
Sales $ 3,800,000
Operating expenses 3,154,000
Net operating income 646,000
Interest and taxes:
Interest expense $ 111,000
Tax expense 198,000 309,000
Net income $ 337,000

The company paid dividends of $245,000 last year. The “Investment in Buisson, S.A.,” on the balance sheet represents an investment in the stock of another company. The company's minimum required rate of return of 15%.

Required:

1. Compute the company's average operating assets for last year.

2. Compute the company’s margin, turnover, and return on investment (ROI) for last year. (Round "Margin", "Turnover" and "ROI" to 2 decimal places.)

3. What was the company’s residual income last year?
Answer #1

Average operating assets

= [(2,541,000-254,000-407,000)+(2,602,000-254,000-428,000)]/2

= 1,900,000

Margin = Net operating income /sales

= 646,000/3,800,000

= 17%

Turnover = Sales /Average operating assets

= 3,800,000/1,900,000

= 2

ROI = Margin*Turnover

= 17*2

= 34%

Residual income = Net operating income - required return

= 646,000 - (1,900,000*15%)

= 361,000